# Discrete conformal structures on surfaces with boundary (II)—Rigidity and Existence

### v1

@@ Appendix A Proofs of Lemmas
For simplicity, we set $h_{t}=h_{rs},\ q_{t}=q_{rs}$ and assume $||v_{i}||=||v_{j}||=||v_{k}||=1$. Please refer to Figure 3.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{i}$, by Proposition 2.1 we have $-v_{i}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{i}\ast c_{ik})$. By Proposition 2.2 $\mathrm{(i)}$, $c_{ik}\ast c_{ijk}=-\cosh h_{j}$. Moreover, by Proposition 2.2 $\mathrm{(ii)}$, if $c_{ik}$ and $v_{i}$ are on same sides of the hyperplane $v_{i}^{\bot}$, then
@@
if $c_{ik}$ and $v_{i}$ are on opposite sides of the hyperplane $v_{i}^{\bot}$, then
@@
Thus $v_{i}\ast c_{ik}=-\sinh d_{ik}$. Similarly, if $c_{ijk}$ and $v_{i}$ are on same sides of the hyperplane $v_{i}^{\bot}$, then $v_{i}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c_{ijk},v_{i}^{\bot}))=+\sinh(-q_{i})=-\sinh q_{i}$; if $c_{ijk}$ and $v_{i}$ are on opposite sides of the hyperplane $v_{i}^{\bot}$, then $v_{i}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{ijk},v_{i}^{\bot}))=-\sinh(+q_{i})=-\sinh q_{i}$. Thus $v_{i}\ast c_{ijk}=-\sinh q_{i}$. Then $\sinh q_{i}=\cosh h_{j}\cdot\sinh d_{ik}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{i}$, we have $-v_{i}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{i}\ast c_{ij})=-\cosh h_{k}\cdot(v_{i}\ast c_{jk})$. Similar arguments can imply $v_{i}\ast c_{ij}=-\sinh d_{ij}$. Then $\sinh q_{i}=\cosh h_{k}\cdot\sinh d_{ij}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ik})=-\cosh q_{j}\cdot(v^{\prime}_{i}\ast c^{\prime}_{ik})$. Similarly, $v^{\prime}_{i}\ast c_{ijk}=-\sinh h_{i}$ and $v^{\prime}_{i}\ast c^{\prime}_{ik}=-\sinh\theta_{ik}$. Then $\sinh h_{i}=\cosh q_{j}\cdot\sinh\theta_{ik}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ij})=-\cosh q_{k}\cdot(v^{\prime}_{i}\ast c^{\prime}_{ij})$. Similarly, $v^{\prime}_{i}\ast c_{ijk}=-\sinh h_{i}$ and $v^{\prime}_{i}\ast c^{\prime}_{ij}=-\sinh\theta_{ij}$. Then $\sinh h_{i}=\cosh q_{k}\cdot\sinh\theta_{ij}$. Q.E.D.
@@
If $c_{ijk}\in\mathbb{R}^{3}$ lies in the open domain $D_{\mathrm{II}}$, then we swap $h$ and $q$, and $\theta$ and $d$.
@@
Suppose $c_{ijk}$ lies in the domain $D_{\mathrm{V}}$. In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{i}$, by Proposition 2.1, we have $-v_{i}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{i}\ast c_{ik})$. By Proposition 2.2 $\mathrm{(iii)}$, $v_{i}\ast c_{ijk}=\cosh(d_{\mathbb{H}}(v_{i}^{\bot},c_{ijk}^{\bot}))=\cosh(-q_{i})=\cosh q_{i}$. Moreover, by Proposition 2.2 $\mathrm{(ii)}$, $v_{i}\ast c_{ik}=+\sinh d_{\mathbb{H}}(v_{i}^{\bot},c_{ik})=+\sinh(-d_{ik})=-\sinh d_{ik}$. If $c_{ik}$ and $c_{ijk}$ are on same sides of the hyperplane $c_{ijk}^{\bot}$, then
@@
If $c_{ik}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then
@@
Thus $c_{ik}\ast c_{ijk}=-\sinh h_{j}$. Then $\cosh q_{i}=-\sinh h_{j}\cdot\sinh d_{ik}$. By Lemma 2.5, it is easy to check that $d_{ik}<0$. Hence $h_{j}>0$. Then $d_{\mathbb{H}}(c_{ik},c_{ijk}^{\bot})=h_{j}$, which implies that $c_{ik}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{i}$, we have $-v_{i}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{i}\ast c_{ij})=-\cosh q_{i}$. Similarly, $v_{i}\ast c_{ij}=-\sinh d_{ij}$. If $c_{ij}$ and $c_{ijk}$ are on same sides of the hyperplane $c_{ijk}^{\bot}$, then $c_{ij}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c_{ij},c_{ijk}^{\bot}))=+\sinh(-h_{k})=-\sinh h_{k}$. If $c_{ij}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c_{ij}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{ij},c_{ijk}^{\bot}))=-\sinh h_{k}$. Thus $c_{ij}\ast c_{ijk}=-\sinh h_{k}$. Then $\cosh q_{i}=-\sinh h_{k}\cdot\sinh d_{ij}$. Similarly, $d_{ij}<0$ by Lemma 2.5 and hence $h_{k}>0$. So $d_{\mathbb{H}}(c_{ij},c_{ijk}^{\bot})=h_{k}$, which implies that $c_{ij}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{j}$, we have $-v_{j}\ast c_{ijk}=(c_{jk}\ast c_{ijk})(v_{j}\ast c_{jk})$. By Proposition 2.2 $\mathrm{(iii)}$, $v_{j}\ast c_{ijk}=-\cosh(d_{\mathbb{H}}(v_{j}^{\bot},c_{ijk}^{\bot}))=-\cosh q_{j}$. Moreover, by Proposition 2.2 $\mathrm{(ii)}$, $c_{jk}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{jk},c_{ijk}^{\bot}))=-\sinh h_{i}$. Similarly, $v_{j}\ast c_{jk}=-\sinh d_{jk}$. Then $\cosh q_{j}=\sinh h_{i}\cdot\sinh d_{jk}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{j}$, we have $-v_{j}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{j}\ast c_{ij})$. Since $v_{j}\ast c_{ijk}=-\cosh q_{j}$, $v_{j}\ast c_{ij}=-\sinh d_{ji}$ and $c_{ij}\ast c_{ijk}=-\sinh h_{k}$, then $\cosh q_{j}=\sinh h_{k}\cdot\sinh d_{ji}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{k}$, we have $-v_{k}\ast c_{ijk}=(c_{jk}\ast c_{ijk})(v_{k}\ast c_{jk})$. Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{jk}=-\sinh d_{kj}$ and $c_{jk}\ast c_{ijk}=-\sinh h_{i}$, then $\cosh q_{k}=\sinh h_{i}\cdot\sinh d_{kj}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{k}$, we have $-v_{k}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{k}\ast c_{ik})$. Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{ik}=-\sinh d_{ki}$ and $c_{ik}\ast c_{ijk}=-\sinh h_{j}$, then $\cosh q_{k}=\sinh h_{j}\cdot\sinh d_{ki}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ik})$. Since $v^{\prime}_{i}\ast c_{ijk}=-\cosh h_{i}$, $v^{\prime}_{i}\ast c^{\prime}_{ik}=-\sinh\theta_{ik}$ and $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$, then $\cosh h_{i}=\sinh q_{j}\cdot\sinh\theta_{ik}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ij})$. Since $v^{\prime}_{i}\ast c_{ijk}=-\cosh h_{i}$, $v^{\prime}_{i}\ast c^{\prime}_{ij}=-\sinh\theta_{ij}$ and $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$, then $\cosh h_{i}=\sinh q_{k}\cdot\sinh\theta_{ij}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{jk}v^{\prime}_{j}$, we have $-v^{\prime}_{j}\ast c_{ijk}=(c^{\prime}_{jk}\ast c_{ijk})(v^{\prime}_{j}\ast c^{\prime}_{jk})$. By Proposition 2.2 $\mathrm{(iii)}$, $v^{\prime}_{j}\ast c_{ijk}=-\cosh(d_{\mathbb{H}}((v^{\prime}_{j})^{\bot},c_{ijk}^{\bot}))=-\cosh h_{j}$. Since $c^{\prime}_{jk}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{jk}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c^{\prime}_{jk},c_{ijk}^{\bot}))=-\sinh(-q_{i})=+\sinh q_{i}$. Due to $v^{\prime}_{j}\ast c^{\prime}_{jk}=-\sinh\theta_{jk}$, then $\cosh h_{j}=-\sinh q_{i}\cdot\sinh\theta_{jk}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{j}$, we have $-v^{\prime}_{j}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{j}\ast c^{\prime}_{ij})$. Note that $v^{\prime}_{j}\ast c_{ijk}=-\cosh h_{j}$, $v^{\prime}_{j}\ast c^{\prime}_{ij}=-\sinh\theta_{ji}$ and $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$, then $\cosh h_{j}=\sinh q_{k}\cdot\sinh\theta_{ji}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{jk}v^{\prime}_{k}$, we have $-v^{\prime}_{k}\ast c_{ijk}=(c^{\prime}_{jk}\ast c_{ijk})(v^{\prime}_{k}\ast c^{\prime}_{jk})$. Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{jk}=-\sinh\theta_{kj}$ and $c^{\prime}_{jk}\ast c_{ijk}=\sinh q_{i}$, then $\cosh h_{k}=-\sinh q_{i}\cdot\sinh\theta_{kj}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{k}$, we have $-v^{\prime}_{k}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{k}\ast c^{\prime}_{ik})$. Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{ik}=-\sinh\theta_{ki}$ and $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$, then $\cosh h_{k}=\sinh q_{j}\cdot\sinh\theta_{ki}$.
@@
Suppose $c_{ijk}$ lies in the domain $D_{\mathrm{II}}$. In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{i}$, we have $-v_{i}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{i}\ast c_{ik})$. By Proposition 2.2 $\mathrm{(iii)}$, $v_{i}\ast c_{ijk}=-\cosh q_{i}$ and $c_{ik}\ast c_{ijk}=-\sinh h_{j}$. Since $v_{i}\ast c_{ik}=-\sinh d_{\mathbb{H}}(v_{i}^{\bot},c_{ik})=-\sinh(+d_{ik})=-\sinh d_{ik}$, then $\cosh q_{i}=\sinh h_{j}\cdot\sinh d_{ik}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{i}$, we have $-v_{i}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{i}\ast c_{ij})=\cosh q_{i}$. Similarly, $v_{i}\ast c_{ij}=-\sinh d_{ij}$ and $c_{ij}\ast c_{ijk}=-\sinh h_{k}$. Then $\cosh q_{i}=\sinh h_{k}\cdot\sinh d_{ij}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{j}$, we have $-v_{j}\ast c_{ijk}=(c_{jk}\ast c_{ijk})(v_{j}\ast c_{jk})$. By Proposition 2.2 $\mathrm{(iii)}$, then $v_{j}\ast c_{ijk}=-\cosh(d_{\mathbb{H}}(v_{j}^{\bot},c_{ijk}^{\bot}))=-\cosh q_{j}$. Since $c_{jk}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c_{jk}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{jk},c_{ijk}^{\bot}))=-\sinh(-h_{i})=+\sinh h_{i}$. Due to $v_{j}\ast c_{jk}=-\sinh d_{jk}$, then $\cosh q_{j}=-\sinh h_{i}\cdot\sinh d_{jk}$.

### v3

@@ Appendix A Proofs of Lemmas
For simplicity, we set $h_{t}=h_{rs}$ and $q_{t}=q_{rs}$, and assume $\|v_{i}\|=\|v_{j}\|=\|v_{k}\|=1$. Please refer to Figure 3.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{i}$, by Proposition 2.1, we have $-v_{i}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{i}\ast c_{ik})$. By Proposition 2.2 (i), $c_{ik}\ast c_{ijk}=-\cosh h_{j}$. Moreover, by Proposition 2.2 (ii), if $c_{ik}$ and $v_{i}$ lie on the same side of the hyperplane $v_{i}^{\bot}$, then
@@
if $c_{ik}$ and $v_{i}$ lie on opposite sides of the hyperplane $v_{i}^{\bot}$, then
@@
In either case, $v_{i}\ast c_{ik}=-\sinh d_{ik}$. Similarly, if $c_{ijk}$ and $v_{i}$ lie on the same side of the hyperplane $v_{i}^{\bot}$, then
$v_{i}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c_{ijk},v_{i}^{\bot}))=+\sinh(-q_{i})=-\sinh q_{i};$
if $c_{ijk}$ and $v_{i}$ lie on opposite sides of the hyperplane $v_{i}^{\bot}$, then
$v_{i}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{ijk},v_{i}^{\bot}))=-\sinh(+q_{i})=-\sinh q_{i}.$
In either case, $v_{i}\ast c_{ijk}=-\sinh q_{i}$. Consequently, $\sinh q_{i}=\cosh h_{j}\cdot\sinh d_{ik}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{i}$, we have $-v_{i}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{i}\ast c_{ij})=-\cosh h_{k}\cdot(v_{i}\ast c_{ij})$. Similar arguments can imply $v_{i}\ast c_{ij}=-\sinh d_{ij}$. Thus $\sinh q_{i}=\cosh h_{k}\cdot\sinh d_{ij}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ik})=-\cosh q_{j}\cdot(v^{\prime}_{i}\ast c^{\prime}_{ik})$. Similarly, $v^{\prime}_{i}\ast c_{ijk}=-\sinh h_{i}$ and $v^{\prime}_{i}\ast c^{\prime}_{ik}=-\sinh\theta_{ik}$. Thus $\sinh h_{i}=\cosh q_{j}\cdot\sinh\theta_{ik}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ij})=-\cosh q_{k}\cdot(v^{\prime}_{i}\ast c^{\prime}_{ij})$. Similarly, $v^{\prime}_{i}\ast c_{ijk}=-\sinh h_{i}$ and $v^{\prime}_{i}\ast c^{\prime}_{ij}=-\sinh\theta_{ij}$. Thus $\sinh h_{i}=\cosh q_{k}\cdot\sinh\theta_{ij}$. Q.E.D.
@@
If $c_{ijk}\in\mathbb{R}^{3}$ lies in the open domain $D_{\mathrm{II}}$, then we swap $h$ with $q$, and $\theta$ with $d$.
@@
Assume $c_{ijk}\in D_{\mathrm{V}}$. In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{i}$, by Proposition 2.1, we have $-v_{i}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{i}\ast c_{ik})$. By Proposition 2.2 (iii), $v_{i}\ast c_{ijk}=\cosh(d_{\mathbb{H}}(v_{i}^{\bot},c_{ijk}^{\bot}))=\cosh(-q_{i})=\cosh q_{i}$. Moreover, by Proposition 2.2 (ii), $v_{i}\ast c_{ik}=+\sinh d_{\mathbb{H}}(v_{i}^{\bot},c_{ik})=+\sinh(-d_{ik})=-\sinh d_{ik}$. If $c_{ik}$ and $c_{ijk}$ lie on the same side of the hyperplane $c_{ijk}^{\bot}$, then
@@
If $c_{ik}$ and $c_{ijk}$ lie on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then
@@
In either case, $c_{ik}\ast c_{ijk}=-\sinh h_{j}$. Consequently, $\cosh q_{i}=-\sinh h_{j}\cdot\sinh d_{ik}$. By Lemma 2.5, it is easy to check that $d_{ik}<0$. It follows that $h_{j}>0$ and hence $d_{\mathbb{H}}(c_{ik},c_{ijk}^{\bot})=h_{j}$. This implies that $c_{ik}$ and $c_{ijk}$ lie on opposite sides of $c_{ijk}^{\bot}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{i}$, we have $-v_{i}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{i}\ast c_{ij})=-\cosh q_{i}$. Similarly, $v_{i}\ast c_{ij}=-\sinh d_{ij}$. If $c_{ij}$ and $c_{ijk}$ lie on the same side of the hyperplane $c_{ijk}^{\bot}$, then $c_{ij}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c_{ij},c_{ijk}^{\bot}))=+\sinh(-h_{k})=-\sinh h_{k}$. If $c_{ij}$ and $c_{ijk}$ lie on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c_{ij}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{ij},c_{ijk}^{\bot}))=-\sinh h_{k}$. In either case, $c_{ij}\ast c_{ijk}=-\sinh h_{k}$. Consequently, $\cosh q_{i}=-\sinh h_{k}\cdot\sinh d_{ij}$. Similarly, $d_{ij}<0$ by Lemma 2.5 and hence $h_{k}>0$. So $d_{\mathbb{H}}(c_{ij},c_{ijk}^{\bot})=h_{k}$, which implies that $c_{ij}$ and $c_{ijk}$ lie on opposite sides of $c_{ijk}^{\bot}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{j}$, we have $-v_{j}\ast c_{ijk}=(c_{jk}\ast c_{ijk})(v_{j}\ast c_{jk})$. By Proposition 2.2 (iii), $v_{j}\ast c_{ijk}=-\cosh(d_{\mathbb{H}}(v_{j}^{\bot},c_{ijk}^{\bot}))=-\cosh q_{j}$. Moreover, by Proposition 2.2 (ii), $c_{jk}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{jk},c_{ijk}^{\bot}))=-\sinh h_{i}$. Similarly, $v_{j}\ast c_{jk}=-\sinh d_{jk}$. Thus $\cosh q_{j}=\sinh h_{i}\cdot\sinh d_{jk}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{j}$, we have $-v_{j}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{j}\ast c_{ij})$. Since $v_{j}\ast c_{ijk}=-\cosh q_{j}$, $v_{j}\ast c_{ij}=-\sinh d_{ji}$, and $c_{ij}\ast c_{ijk}=-\sinh h_{k}$, it follows that $\cosh q_{j}=\sinh h_{k}\cdot\sinh d_{ji}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{k}$, we have $-v_{k}\ast c_{ijk}=(c_{jk}\ast c_{ijk})(v_{k}\ast c_{jk})$. Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{jk}=-\sinh d_{kj}$, and $c_{jk}\ast c_{ijk}=-\sinh h_{i}$, it follows that $\cosh q_{k}=\sinh h_{i}\cdot\sinh d_{kj}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{k}$, we have $-v_{k}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{k}\ast c_{ik})$. Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{ik}=-\sinh d_{ki}$, and $c_{ik}\ast c_{ijk}=-\sinh h_{j}$, it follows that $\cosh q_{k}=\sinh h_{j}\cdot\sinh d_{ki}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ik})$. Since $v^{\prime}_{i}\ast c_{ijk}=-\cosh h_{i}$, $v^{\prime}_{i}\ast c^{\prime}_{ik}=-\sinh\theta_{ik}$, and $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$, it follows that $\cosh h_{i}=\sinh q_{j}\cdot\sinh\theta_{ik}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ij})$. Since $v^{\prime}_{i}\ast c_{ijk}=-\cosh h_{i}$, $v^{\prime}_{i}\ast c^{\prime}_{ij}=-\sinh\theta_{ij}$, and $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$, it follows that $\cosh h_{i}=\sinh q_{k}\cdot\sinh\theta_{ij}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{jk}v^{\prime}_{j}$, we have $-v^{\prime}_{j}\ast c_{ijk}=(c^{\prime}_{jk}\ast c_{ijk})(v^{\prime}_{j}\ast c^{\prime}_{jk})$. By Proposition 2.2 (iii), $v^{\prime}_{j}\ast c_{ijk}=-\cosh(d_{\mathbb{H}}((v^{\prime}_{j})^{\bot},c_{ijk}^{\bot}))=-\cosh h_{j}$. Since $c^{\prime}_{jk}$ and $c_{ijk}$ lie on opposite sides of the hyperplane $c_{ijk}^{\bot}$, it follows that $c^{\prime}_{jk}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c^{\prime}_{jk},c_{ijk}^{\bot}))=-\sinh(-q_{i})=+\sinh q_{i}$. Due to $v^{\prime}_{j}\ast c^{\prime}_{jk}=-\sinh\theta_{jk}$, we conclude that $\cosh h_{j}=-\sinh q_{i}\cdot\sinh\theta_{jk}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{j}$, we have $-v^{\prime}_{j}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{j}\ast c^{\prime}_{ij})$. Since $v^{\prime}_{j}\ast c_{ijk}=-\cosh h_{j}$, $v^{\prime}_{j}\ast c^{\prime}_{ij}=-\sinh\theta_{ji}$, and $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$, it follows that $\cosh h_{j}=\sinh q_{k}\cdot\sinh\theta_{ji}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{jk}v^{\prime}_{k}$, we have $-v^{\prime}_{k}\ast c_{ijk}=(c^{\prime}_{jk}\ast c_{ijk})(v^{\prime}_{k}\ast c^{\prime}_{jk})$. Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{jk}=-\sinh\theta_{kj}$, and $c^{\prime}_{jk}\ast c_{ijk}=\sinh q_{i}$, it follows that $\cosh h_{k}=-\sinh q_{i}\cdot\sinh\theta_{kj}$.
[…]
@@
Assume $c_{ijk}\in D_{\mathrm{II}}$. In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{i}$, we have $-v_{i}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{i}\ast c_{ik})$. By Proposition 2.2 (iii), $v_{i}\ast c_{ijk}=-\cosh q_{i}$ and $c_{ik}\ast c_{ijk}=-\sinh h_{j}$. Since $v_{i}\ast c_{ik}=-\sinh d_{\mathbb{H}}(v_{i}^{\bot},c_{ik})=-\sinh(+d_{ik})=-\sinh d_{ik}$, it follows that $\cosh q_{i}=\sinh h_{j}\cdot\sinh d_{ik}$.